Calculate the frequency of the red light emitted by a neon sign with a wavelength of 662 nm
Answer #1

We know that,

c = n * l

Where c is speed of light with value 3*10^8 m/s , n is frequency and l is wavelength.

l = 662 nm = 662*10^-9 m

n = c/l = 3*10^8 / 662*10^-9 = 4.532*10^14 Hz ...Answer